OpenStudy (anonymous):

I cannot figure this out for the life of me ;( Solve the following system of equations and show all work. y = x^2 + 3 y = x + 5

OpenStudy (radar):

Note that the right hand side of each equation is equal to y......soooo they must also be equal to each other. Agree ?

OpenStudy (radar):

We can now write: x^2 +3 = X + 5 which can also be written as: x^2 - x - 2 = 0 Do you agree? Or have you just went to the "looking around mode?"

OpenStudy (radar):

Factor and solve.

OpenStudy (radar):

(x-2)(x+1) = 0 x=2 x=-1 Now solve for y using the values found for x y= x +5 y=-1 +5=4 y=2 + 5 =7 Now insert in first equation: y = x^2 + 3 y= 2^2 + 3 = 7 y= (-1)^2 + 3 = 4 (2,7), (-1,4)

OpenStudy (radar):

Not a life saving problem.